Calculate molality of 10gm of glucose in? 90gm of water

Grade:12th

1 Answers

Arun
6 years ago
In order to know your answer firstly know about molality
Molality may be defined as the number of moles of solute that are  dissloved in a kg of water .
so firstly we have to calculate the moles of glucose
10*1/180.16 = 0.0555 ( as the molar mass of glucose= 180.1559)
These are dissloved in 90 g of water.
first convert mass into kilogram =90/1000= 0.090 kg
0.090 kg water = 0.0555 moles of glucose
so 1 kg of water=1*0.0555/0.090 = 0.6166 moles of glucose
round about of 0.6166 will be 0.617 which is our answer